### v1

# On the handling of method failure in comparison studies

Abstract
Comparison studies in methodological research are intended to compare methods in an evidence-based manner, offering guidance to data analysts to select a suitable method for their application. To provide trustworthy evidence, they must be carefully designed, implemented, and reported, especially given the many decisions made in planning and running. A common challenge in comparison studies is to handle the “failure” of one or more methods to produce a result for some (real or simulated) data sets, such that their performances cannot be measured in those instances. Despite an increasing emphasis on this topic in recent literature (focusing on non-convergence as a common manifestation), there is little guidance on proper handling and interpretation, and reporting of the chosen approach is often neglected. This paper aims to fill this gap and provides practical guidance for handling method failure in comparison studies. In particular, we show that the popular approaches of discarding data sets yielding failure (either for all or the failing methods only) and imputing are inappropriate in most cases. We also discuss how method failure in published comparison studies—in various contexts from classical statistics and predictive modeling—may manifest differently, but is often caused by a complex interplay of several aspects. Building on this, we provide recommendations derived from realistic considerations on suitable fallbacks when encountering method failure, hence avoiding the need for discarding data sets or imputation. Finally, we illustrate our recommendations and the dangers of inadequate handling of method failure through two illustrative comparison studies.
1 Introduction
In methodological research, comparison studies aim to compare the performance of methods to provide empirical evidence on their behavior and help data analysts choose appropriate methods for their setting. To be reliable, they require careful planning of the design, execution, and reporting of the results. These issues have gained increasing attention in recent years, both in the context of simulation studies based on artificially generated data (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024) and in the context of real data-based benchmarking studies (Boulesteix et al., , 2017; Friedrich and Friede, , 2024).
In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure”. Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study. The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024), often focusing on non-convergence. However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form. This is illustrated by our review of published comparison studies from classical statistics and predictive modeling in Section 2 which shows that (reported) handlings often differ greatly between comparison studies. Nevertheless, these studies are positive examples of the reporting and the handling of method failure as this issue often goes unreported in many other comparison studies. For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur” (Hinds et al., , 2018). Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) reported convergence as a performance measure. This contrasts with our perception—both through our own experiences and informal discussions with colleagues—that the problem of method failure is highly prevalent. In particular, signs of method failure are often traceable in the code provided for reproducibility even when not reported in the manuscript. This is often clear through the use of R functions try() or tryCatch() for the handling of errors, for example.
To illustrate the difficulty of handling method failure adequately, imagine conducting a comparison study in statistics or predictive modeling to compare the performance of a selection of methods. In the context of classical statistics, you likely focus on comparing the methods’ performances in “estimating one or more population quantities” (Morris et al., , 2019), sometimes called “estimand(s)” and base your analysis on data simulated using a certain data-generating mechanism (DGM). In predictive modeling, you are likely to compare the methods’ predictive performances, that is, their ability to make predictions on a new set of observations. Here, the analysis is typically performed on real benchmarking data sets. For simplicity, we refer to both settings as “simulation studies” and “benchmark studies”, respectively (though admitting that there may be intermediate situations, e.g. simulation studies investigating predictive performance on artificial data).
During your experiments, you encounter the failure of a method in one or more of the following ways: (i) you receive an error message or output of NA (“not applicable”) or NaN (“not a number”), implying that a certain calculation could not be completed (e.g., non-convergence), (ii) your system crashes, or (iii) the experiment runs for an excessively long time, forcing you to abort.
Table 1: Overview table of (a) a fictive simulation study and (b) a fictive benchmark study encountering method failure. This results in undefined estimates values for the simulation study (a) and undefined performance values for the benchmark study (b), marked in blue. In both cases, this complicates the aggregation of performance of the given method.
(a) Fictive simulation study: overview of estimates produced by three methods under investigation (columns) across $4$ repetitions, i.e. simulated data sets (rows), in one simulation scenario, where the true value is equal to $4$. An estimates value is obtained for each of the three methods and four simulation repetitions. Bias is commonly calculated as the mean deviation of the estimates from the true value across the $4$ repetitions. Method failure first leads to an undefined estimates value in the given repetition, which then complicates the estimation of performance of the affected method.
Repetition Method 1 Method 2 Method 3
1 3.89 4.23 4.08
2 3.78 4.13 4.11
3 NA 3.69 4.23
4 3.75 4.24 NA
Bias ? 0.07 ?
(b) Fictive benchmark study in predictive modeling: an estimated accuracy is obtained for each of the four data sets (rows) and three methods under investigation (columns). Method failure directly leads to an undefined performance (i.e., accuracy) value for the given method in the given row of the table, complicating the aggregation of its performance over all rows required to obtain an overall performance estimation.
Benchmark data set Method 1 Method 2 Method 3
1 0.85 0.88 0.87
2 0.9 0.91 0.86
3 NA 0.80 0.82
4 0.78 0.76 NA
Average accuracy ? 0.84 ?
Table 1 depicts a hypothetical summary of the results of your study in each simulation repetition/benchmark data set, denoting the undefined values by a blue “NA”. You may now wonder how to accurately determine the methods’ performances from this incomplete information. Note that Table 1 first of all highlights an important distinction that must be made between simulation and benchmark studies regarding the consequences of method failure. In a simulation study focusing on assessing the methods’ performances in estimating an estimand, a method typically produces a single estimate for the entire data set in the given repetition (i.e., row in Table 1(a)). Its performance is estimated as an aggregate over all rows in Table 1(a). In this setting, method failure first leads to an undefined estimate value which then complicates the assessment of the affected method’s performance. In a benchmark study focusing on the predictive performance of methods, on the other hand, a method makes a prediction for each row of the corresponding test data for a given benchmark data set (based on the decision rule generated on the training data). The method’s performance (e.g., accuracy) is then measured directly in each real data set (i.e., each row in Table 1(b)) and its overall performance is averaged over all rows of Table 1(b). Method failure results directly in an undefined performance value.
Looking at Table 1, it is tempting to view method failure as missing data: in each case, the row summarising results contains “holes” where results were expected. We caution against this. Missing values are data that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin, , 2004; Carpenter et al., , 2023). In our context, although we intended to record a value, method failure means that the value does not exist for that particular data set. This is similar to data truncation by death as opposed to the inability to collect data due to the non-appearance of a patient at the doctor’s office (missing data). As a result, some of the most commonly used approaches to handling method failure such as imputation, being “missing data inspired”, are inadequate in most cases. Furthermore, only few are aware that the typical ways and reportings of discarding data sets associated with failure (either for all or the failing methods only) do not enable meaningful method comparisons. We illustrate this in detail in Sections 3 and 4.
In the subsequent part of this paper, we use the term “undefined” instead of “missing” to emphasize the important distinction between method failure and the “regular” missing data context. The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
The rest of this paper is structured as follows. In Section 2, we review and summarize the different manifestations of method failure in published comparison studies. In particular, we discuss how at the core, they often result from a complex interplay of multiple factors, and that an adequate handling requires thorough investigation on a case-by-case basis. Building on this, Section 3 is a compilation of practical recommendations for handling method failure which differentiates between different scopes of the study and research stages in which it may take place. In particular, we promote using fallback strategies when encountering method failure. This approach, if based on thorough and realistic considerations, enables the aggregation of performance across data sets even when method failure occurs. However, considering possible time constraints, we also discuss “slimmed-down” ways in which authors can still implement our recommendations. Finally, in Section 4, we demonstrate our recommendations using two fictive comparison studies that illustrate the complexity of the discussed issues.
2 Method failure in the literature
In our own experience, method failure occurs frequently when conducting comparison studies in statistics and predictive modeling. As already described, a look at the literature suggests that others also encounter this problem, either because method failure is explicitly mentioned in the text or because signs of method failure can be found in the code provided by the authors for reproducibility. Section 2.1 briefly outlines how method failure manifests and is handled in a selection of published studies from the classical statistics and predictive modeling literature. In Section 2.2, we argue that at the core, method failure is a complex interplay of multiple factors, deserving particular attention. Finally, Section 2.3 briefly categorizes different procedures for handling method failure used in these studies and provides a first outlook on the recommendations provided in Section 3.
2.1 Reporting and handling of method failure from the literature: vignettes
The following comparison studies should be viewed as positive examples, being among the few that report the occurrence of method failure, often alongside reasons and strategies for handling the resulting undefined values when summarizing and interpreting the results. The strategies adopted in these studies inspire—but do not completely overlap with—our recommendations given in Section 3. In these cases, we do not intend to criticize the applied approaches but aim to further improve the handling and reporting of method failure for future work.
van Smeden et al., (2016) investigate why different simulation studies offer conflicting minimal “events per variable” recommendations for logistic regression. They particularly observe that separated data sets (i.e., outcome variable is perfectly predicted by one or a combination of multiple covariates) are commonly omitted from the analysis since they lead to convergence issues of the likelihood maximization process. However, van Smeden et al., (2016) demonstrate that the removal of these data sets, in which the association between the covariates and outcomes is particularly strong, substantially affects bias, mean squared error (MSE), and confidence interval width, even if the frequency of separated data sets is small.
In Voeltz, (2024), a simulation study is conducted to compare three frequentist methods for the meta-analysis of diagnostic accuracy studies regarding sensitivity and specificity. Two of the three methods under investigation are occasionally affected by method failure due to non-convergence (while the third one does not employ an iterative process). In the case of non-convergence, one of the two does not return an error but outputs the current, non-converged estimate. As these outputs are often used in practice nevertheless, the authors also include them in the performance comparison. For the remaining method affected by non-convergence, they discard the data sets for which it fails and use the proportion of simulation repetitions in which the methods converge as an additional performance measure.
An extreme example of excluding performance data is to remove an entire method. This is done by Ruxton and Neuhäuser, (2013) in a simulation study to compare the empirical coverage of methods commonly used to calculate confidence intervals for the odds ratio. Given a binary outcome and binary exposure variable, some methods fail due to division by 00 if the corresponding $2\times 2$ contingency table contains one or more zero-entries.
Masaoud and Stryhn, (2010) compare the performance of different marginal and random effects estimation methods for the analysis of binary longitudinal data (e.g., GEE, Bayesian Markov Chain Monte Carlo) in a simulation study. With data simulated using two different data-generating mechanisms (i.e., stemming from marginal and random effects models), methods are compared regarding bias and coverage of the confidence intervals (as well as efficiency for data simulated from the marginal model). The authors encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010), for which the handling varies for the different methods affected. They either (i) change the software from R to SAS, (ii) switch the optimization technique used internally by the method, or (iii) for GEE, resort to using model-based standard errors if the method fails to produce meaningful robust standard errors. However, if method failure cannot be solved using these measures, they discard the corresponding data set.
Dunias et al., (2024) conduct a simulation study to compare the performance of four hyperparameter tuning procedures (including $k$-fold cross validation and the one-standard error rule for tuning) for prediction methods developed using predictive modeling methods such as Ridge, LASSO, or Random Forest. For some data sets, LASSO does not select any predictor variable, leading to a constant predictor. Resulting in the inability to assess the performance in terms of the calibration slope, the “[missing] calibration slopes [are] replaced by the maximum value of calibration slope for the corresponding scenario” (Dunias et al., , 2024). Note that this corresponds to imputation by the worst “observable” value since higher values of the calibration slope generally indicate that “risk predictions are too narrowly distributed” (Dunias et al., , 2024).
An approach to handling method failure by imputation is applied by Bischl et al., (2013) in their comparison study of $11$ methods for classification on $49$ real and simulated (benchmarking) data sets. They use misclassification error as performance measure combined with $100$ subsampling (also called “Monte Carlo cross-validation”) repetitions. When encountering method failure, they impute based on the proportion of repetitions in which the affected method fails. If below $20\%$, they impute by sampling from a normal distribution that is estimated from the remaining performance values of that model. This way of imputing serves to adhere to the assumptions of the model used to test the generalization performance. If a method fails for more than $20\%$ of repetitions, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al., , 2013).
Herrmann et al., (2021) perform a benchmark study to compare the predictive performance of survival time prediction methods based on high-dimensional multi-omics data using real cancer data sets. They encounter method failure in several cross-validation folds for several methods (including one method with excessive runtimes). Inspired by Bischl et al., (2013), they impute undefined performance values based on the proportion of “failed” cross-validation folds. If it is above $20\%$, they impute by the “worst possible value” of the corresponding metric, else, they impute by the mean performance of the method in the remaining “successful” folds. Note that the latter is equivalent to discarding those data sets in the performance assessment for which the given method fails. In a subsequent project, Nießl et al., (2022) investigate the effect of different imputation procedures on the performance assessment and ranking of the methods. They consider alternative procedures including imputation by the worst possible value and by a value that weights the method’s overall proportion of undefined performance values and its performance in the remaining folds. They show that the approach to imputing the undefined performance values has a notable effect on the performance ranking between the methods.
Another example can be found in the work of Hornung et al., (2024). They assess the performance of prediction methods that can handle block-wise multi-omics data by artificially inducing specific patterns of block-wise missing values into initially complete multi-omics data. Three of the considered methods fail for some data sets (with a maximum proportion of $34\%$). In some rare cases, failure is caused by memory issues. More commonly, however, the methods are “not applicable” (Hornung et al., , 2024) if the training and test data sets differ in the induced pattern of missing data blocks in certain ways. The resulting undefined performance values are addressed by discarding the corresponding data sets for all methods with the reasoning that discarding them only for the failing methods would lead to confounding by the induced pattern of block-wise missing values. Note that Hornung et al., (2024) argue that instances of memory issues are likely caused by a specific missingness pattern induced into the training data.
Fernández-Delgado et al., (2014) perform an extensive comparison study of supervised classification methods (“classifiers”) from several model families. The classifiers are assessed in $121$ (real) benchmarking data sets in combination with $4$-fold cross-validation and “accuracy” as performance measure (which is then averaged over all benchmarking data sets). Additionally, they compute the Friedman ranking (Sheskin, , 2011) based on accuracy. They encounter method failure in $2.1\%$ of classifier-data set combinations. As reasons for failure, they list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014). Note that equal predictor values of all observations in the training data lead to the failure of the corresponding method since this violates its requirements. Memory issues also lead to method failure with large data sets, while particularly low numbers of observations in some classes in small data sets lead to errors for some classifiers. The authors discard the corresponding data sets for the failing methods to obtain each classifier’s accuracy. For the Friedman ranking, which requires an identical number of accuracy values for all classifiers, they impute undefined accuracy values with the mean value across all classifiers for this given data set.
Gijsbers et al., (2024) introduce a benchmarking tool for automated machine learning (“AutoML”) frameworks, which are considered as the “methods” under comparison. They conduct a comparison study of nine such methods based on various data sets using the Area Under the Curve (AUC) as performance metric. They categorize the reasons for failure into (i) exceeding available memory (or other memory-related issues), (ii) exceeding time limits, (iii) failure related to data set characteristics (such as highly imbalanced data), and (iv) failure caused by errors in the implementation of the AutoML framework. Categories (i) and (ii) are reported as the most common. Gijsbers et al., (2024) argue against ignoring undefined performance values, stating that they do not occur at random but are correlated with data set characteristics. They impute by the performance of a constant predictor.
2.2 Typical manifestations of method failure
Summarizing from the previous Section 2.1, there are common ways in which method failure often reveals itself to authors of comparison studies; see Table 2 for an additional overview.
Table 2: Common manifestations of method failure
Manifestation Explanation Examples from literature
Non-convergence
and other
calculation issues
Method cannot perform required calculations (within maximum number of iterations), therefore fails to produce a stable output. Cannot be solved by increasing memory or time budgets.
$\rightarrow$ Example: - division by zero - attempt to invert singular matrix $\rightarrow$ Can result in: - error message - non-meaningful output such as NA, NaN, output of $\infty$, or inflated estimation (for estimation task)
van Smeden et al., (2016)
Fernández-Delgado et al., (2014)
Memory
issues
Method consumes more memory than available to perform required computations.
$\rightarrow$ Can result in: system crashes (“fatal error”)
Hornung et al., (2024)
Fernández-Delgado et al., (2014)
Runtime
issues
Method fails to perform required computations within pre-specified or feasible time limits.
$\rightarrow$ “Failure by design” of study
$\rightarrow$ Real users would likely obtain valid output
Herrmann et al., (2021)
Gijsbers et al., (2024)
Non-convergence and other calculation issues  Authors often encounter the inability of a method to perform a calculation necessary for prediction or estimation, either implied through an error, a warning message, or a method output that is not meaningful. For methods employing an iterative procedure (e.g., an iterative maximum-likelihood procedure), this is often referred to as “non-convergence”, i.e., the method fails to produce a stable and valid output in the pre-specified number of iterations.
Memory issues  Even before exceeding possible time budgets, some methods crash due to memory issues. That is, the given method consumes more memory than is available. Gijsbers et al., (2024) additionally list segmentation faults as memory-related errors.
Runtime issues  A variant of method failure manifested and reported in existing comparison studies is excessive runtime. Note that in some comparison studies, fixed time budgets are predefined (possibly with an additional leniency period) such that exceeding them is declared method failure (Gijsbers et al., , 2024). Other studies do not predefine fixed budgets but find that, due to excessive runtimes, running the entire experiment for the affected method is impossible and must therefore be aborted (Herrmann et al., , 2021). Note that in contrast to the other two manifestations, method failure is actively declared by the authors when runtime issues occur. That is, running the methods under investigation on hundreds or thousands of data sets in the study creates the necessity of time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output. In particular, in many cases, a corresponding real-world user, running the method on a single data set only, would likely obtain a valid output. This has important implications for the handling of method failure, which we will discuss in Section 3.2.
At first glance, it may seem like calculation, memory, and runtime issues are separate reasons for method failure, such that authors report specific “reasons” for failure for different methods within their study or even different “reasons” for a single method in different instances. For instance, some authors report non-convergence as a “reason” for failure in one method–data set combination and memory issues in another. However, upon closer examination, it becomes evident that these are not so much reasons but rather manifestations of method failure that are just the tip of the iceberg. In essence, method failure (or what might appear as such) is often the result of an individual and complex interplay of many different factors.
For example, Kriegel et al., (2017) show that runtimes can depend on numerous factors: “You might be surprised by the amount of differences in runtime that we point out, spanning up to several orders of magnitude in different implementations of one and the same algorithm. We found major performance differences even within the same toolkit, in different versions or releases: due to bugs, algorithmic improvement (having an indirect impact on the tested algorithm, e.g., in data handling), regressions, and environment changes. Finally, it is also noteworthy that good or bad choice of parameters for different algorithms can lead to void runtime comparisons”.
It is important to emphasize that differences in hardware (e.g., available memory) and/or software (e.g., operating system) can also affect assessment. As an extreme example of software dependency, consider that Herrmann et al., (2021) reported that one of the methods under investigation (sparse group lasso) was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu. As an extreme example of hardware dependency, consider that the feasibility of deep learning to train large artificial neural networks tremendously increases with the use of graphics processing units (GPUs). In essence, this means it is important to keep in mind that a computation that fails or is infeasible in terms of runtime and budget on one specific system setup, may work on a different setup. Note that the reverse situation that a computation that can be successfully conducted on one specific system cannot be conducted—i.e., reproduced—using a (slightly) different setup (e.g., during review), is very common and addressed by the term computational reproducibility.
In addition to these (combinations of) factors associated with a method (or its implementation) and the technical issues mentioned above, failure is typically also related to certain characteristics of the given data. That is, some characteristics can complicate required calculations for a method such that they are completely impossible or cannot be completed within the maximum number of iterations (non-convergence), or they exceed available time or memory budgets. For instance, van Smeden et al., (2016) report non-convergence of logistic regression algorithms when the outcome variable is perfectly separated by one or more predictor variables.
To sum up, oftentimes, certain characteristics of the data cause a certain method, in its current implementation, to fail in a certain computational environment.
2.3 Common approaches to handling method failure
The approaches to handling method failure from the published comparison studies in Section 2.1 can generally be divided into two categories. First, some authors resolve method failure so that no undefined values remain. This includes entirely removing failing methods from the study or modifying its implementation (e.g., adjusting some parameters) to make it work for all data sets. Second, other authors summarize the affected method(s) performance(s) despite undefined performance values. Most commonly, the corresponding data sets are discarded either for all or only for the failing methods. Other authors impute undefined performance values. Again, this is often done in various ways, for example, imputation through the worst possible value (when such a value exists) or through the worst (or mean) value achieved by other methods on the particular data set. Additionally, some authors also report the proportion of (real or simulated) data sets for which each method fails.
In the following Section 3, we discuss why imputation and discarding data (either for all or the failing methods only) are inadequate in most cases of method failure. In particular, we recommend alternative and more appropriate approaches that are based on realistic considerations regarding the research stage in which the study takes place and the corresponding targeted users.
3 Principles and recommendations for handling method failure
3.1 Absolute, relative, unconditional and conditional performance
There is no one-size-fits-all solution to handling method failure, exemplified by the examples in Section 2. This section discusses some principles to guide decisions. To illustrate ideas, we begin with two simple examples and some useful terminology. Suppose a simulation study aims to compare the performance of method A with method B in terms of bias. Method A fails in $20\%$ of repetitions while method B does not fail in any. Alternatively, suppose a benchmark study aims to compare the average accuracy of two methods. Again, method A fails in $20\%$ of benchmark data sets while method B does not fail in any. The hypothetical results for the two cases are given in Table 3(b) (a) and (b).
Table 3: Overview table of an illustrative comparison study comparing the performance of (a) two estimation methods regarding bias and (b) two prediction methods regarding average accuracy. Method A fails in $20\%$ of (a) simulation repetitions and (b) benchmark data sets.
(a) Hypothetical simulation study
Conditional
Bias Unconditional (repetitions where
(all repetitions) A does not fail)
Method A - 0.09
Method B 0.14 0.09
(b) Hypothetical benchmark study
Conditional
Average Unconditional (data sets where
accuracy (all data sets) A does not fail)
Method A - 0.85
Method B 0.72 0.85
We contrast the terms unconditional vs. conditional performance and absolute vs. relative performance. Unconditional performance simply means “in all data sets”, while conditional means “in data sets with no method failure”. The unconditional bias of method B is 0.14, while the unconditional bias of method A is not defined. The conditional bias of both methods is 0.09. Parallel descriptions apply to the hypothetical benchmark study in Table 3(b) (b). Absolute performance means “considered on its own” and relative performance means “when compared with another method”. In both Tables 3(b) (a) and 3(b) (b) the relative unconditional performance of methods cannot be quantified because the unconditional performance of method A is undefined. The relative conditional performance can be quantified: in Table 3(b) (a), the two methods have the same bias conditional on non-failure of method A; in Table 3(b) (b), the two methods have the same accuracy conditional on non-failure of method A.
These terms help to distill the problems with method failure in comparison studies. Firstly, the absolute conditional performance of method B should be of little interest. In deciding to use method B, why should we care about its performance when method A, a method we are not going to use, does not fail? We should care about a method’s unconditional performance. The problem is that we cannot compare the methods in terms of their unconditional performance. Comparisons should be performed vertically, or we are not comparing like with like: if we compared the unconditional performance of method B with the conditional performance of method A (which corresponds to discarding data sets for the failing methods), this would reflect poorly on method B, even though both perform equally well for the data sets without failure. This is all the more problematic since method B, in this case, is actually preferable to method A, as it outputs a result for all data sets!
The methods’ conditional performance (“discard corresponding data sets for all methods”) does compare like-with-like. In conducting this comparison, method B’s conditional performance is in both cases (a) and (b) better than its unconditional performance.
3.2 Caution with discarding data and imputation
It is clear that what we would really like to quantify, but cannot, is relative unconditional performance. The superficial solution often proposed when looking at results such as those in Table 3(b) is to consider undefined values due to method failure as missing values and use missing data methods, such as imputation to infer the unconditional performance of method A. After all, results given by different methods tend to be positively correlated, so results given by method B could inform the imputation.
We view it as mistaken to regard method failure as a form of missing data and have been careful to refer to “undefined” rather than “missing” values. The missing data literature defines missing values as those that exist but whose values were not recorded (Carpenter et al., , 2023; Rubin, , 2004). Method failure is far closer to what is called a truncating event in the estimands literature (Kurland et al., , 2009; Kahan et al., , 2024), which denotes a value that does not exist. That is, when a method fails, the result we wished to record simply does not exist for that data set. For this reason, we should be cautious about applying missing data approaches to method failure. In particular, imputing undefined performance values or discarding data sets associated with method failure (either for all methods or the failing methods only) do not enable the comparison of relative unconditional performance. There are two exceptions to this.
First is when a “failure” is recorded but a result could have been obtained; for example, if failure occurs due to excessive runtime but the definition of excessive in the comparison study can safely be judged as shorter than in practice. For cases like this, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”). As this can be seen as missingness mechanism “missing completely at random” (missingness is “unrelated to the observed and unobserved data” (Carpenter et al., , 2023)), discarding the data sets for which the method “fails” or imputing based on the performance results from other methods or data sets may be appropriate. To ensure missingness completely at random, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes. Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets before conducting the study (see Section 3.3.7).
A second exception is when failure occurs for no reason, such that it is safe to compare the unconditional performance of method B with the conditional performance of method A. However, because method failure always happens for a reason – the interplay of a particular method with a particular data set – we do not believe this would ever occur.
An often more suitable approach to handling method failure than discarding data sets or imputing, which we term the “fallback strategy”, involves specifying a backup method when a failure occurs. This strategy is described in more detail in Section 3.3.5.
Table 4: Recommendations on handling method failure
Caution with Explanation
Imputation In most cases: method failure does not lead to “missing data”
$\rightarrow$ Possible exception: Runtime issues
Discarding data
(for all methods or the
failing methods only)
Usually does not provide assessment and comparison of relative unconditional performance
$\rightarrow$ Possible exception: Runtime issues
Alternative recommendations Explanation
Pinpoint interplay causing failure as the basis of any of the following considerations…
Check for method misuse Is the failing method truly developed for given estimation or prediction task?
Check if data matches scope Are targeted users likely to encounter this kind of data and failure?
Be realistic when modifying method implementations Consider changes in implementations to solve method failure that are realistic for targeted users and research stage. Possibly combine with fallback strategies (see below).
Incorporate fallback strategy Incorporate method pipelines “method A if it succeeds in producing an output, method B otherwise”.
$\rightarrow$ Enables assessment of relative unconditional performance.
Report transparently … - failure proportions, ideally as performance measure. - interplay causing failure with sufficient detail: “Method Z, with an overall failure proportion of $X\%$, fails when … and when … .” - chosen handling(s) and justify choices based on above considerations.
Slimmed-down recommendations
Analyze and report simultaneously:
(i) results after discarding data sets for all methods.
(ii) results after discarding data sets for the failing methods.
(iii) proportions of method failure for all methods.
$\rightarrow$ Report that this does not provide assessment of relative unconditional performance.
3.3 Practical recommendations in the presence of method failure
In the following, we present practical recommendations in the presence of method failure that may be more appropriate than imputing or discarding data sets. We advise implementing these recommendations as comprehensively as possible as they contribute to increased usability of comparison studies to users of the methods. However, we are aware that methodological researchers facing time constraints and page limits may be prevented from following them completely. In Section 3.4, we therefore discuss how a “slimmed-down” version of our recommendations can be applied nevertheless.
3.3.1 Pinpoint interplay of method and data characteristic causing failure
We recommend authors of comparison studies investigate and understand for which types of data sets and why a corresponding method fails, beyond the initially perceived symptom(s). Ideally, they should identify if particular (and if different) data characteristics cause the failures, leading to evaluations such as “method A fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”. Having identified the dynamic between the given method and data that causes method failure, authors should base their further handling on the considerations and recommendations provided in the following Sections 3.3.23.3.7.
3.3.2 Check for method misuse
When encountering method failure, authors should first check whether the method was truly developed for the given estimation or prediction task, and the type of data (based on the investigations from the previous Section 3.3.1), and if it is used as intended by the developers. If not the case, what appears initially as method failure may actually be method misuse. Method misuse should not be attributed to the method but to the user, that is, the method should not be penalized for failing on a data set or a task it was not intended for. To avoid method misuse and any associated failure in the first place, authors of comparison studies should familiarize themselves thoroughly with the methods under investigation and read the corresponding user manuals carefully. It may also make sense to consult the developers of the individual methods (and possibly involve them in the study).
Note that there may also be cases where a method is commonly misused by real-world users, especially when they are inexperienced. Depending on which kind of user the study aims at (e.g., inexperienced versus experienced), the occurrence of this kind of failure may be relevant to them and should therefore be included in the study. In these cases, the misuse should be clearly emphasized.
3.3.3 Check if data matches scope of study
If method failure remains after solving or ruling out method misuse, authors should assess if real-world users are likely to encounter such data (and therefore, this method failure) in practice. These considerations must include the scope of the study and are of course not purely objective. Importantly, they should be reported in the study, since they carry valuable information and affect the interpretation of the study’s results.
Consider the case of method failure caused by the method’s inability to make required calculations due to an “extreme” structure of the given simulated data set, for example highly imbalanced classes in the outcome or predictor variable(s), as in White et al., (2024). Depending on the scope of the study, these extreme data sets may be of interest or not. If they are deemed unrealistic for the considered scope and thus not relevant to the readers of the study, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic given the aim of their study, in line with White et al., (2024). Similarly, “too extreme” data may also affect real data based studies. If one of the (real) benchmarking data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely. See Section 4.2 for an illustration.
However, if the data sets causing failure are deemed relevant to the study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation. In particular, they may modify certain parameters of the concerned method(s) (see Section 3.3.4) or resort to a fallback (see Section 3.3.5).
3.3.4 Be realistic when modifying method implementations
As already touched upon, in some instances of method failure, a way of mimicking the behavior of real-world users may be to resort to an alternative implementation of the affected method (see Masaoud and Stryhn, (2010) for an example). This way, method failure may be resolved completely, allowing for the assessment of unconditional (absolute and relative) performance of the methods under investigation.
For instance, one could modify a parameter of the method, such as increasing the maximum number of iterations in the iterative maximum likelihood process or changing the internally employed optimization algorithm. However, authors of comparison studies should not tweak the implementation of the method thoughtlessly. Instead, they should always consider if the applied change in the parameter setting fits the given research stage and corresponding scope of the study. For instance, if the study goal is to understand the methods’ behaviors quite generally (and in earlier phases of research according to the terminology of Heinze et al., (2024)), authors may modify a greater variety of parameters to see when the method is successful and when it fails. However, if the study takes place in later phases and aims to guide inexperienced users, the applied parameter changes should be more restricted to what a naive user of the method would be able and likely to do. In any case, the performance assessment of the given method is only meaningful if the authors stress and justify their parameter choices when reporting the study.
Furthermore, it is important to make sure that the method comparison is still fair, meaning that the effort spent on each method under consideration should not differ significantly between the methods, especially in late phase studies, where neutrality is of crucial importance. Finally, one should be careful about increasing the runtimes of the affected method(s) when running a comparison study. Especially when the number of iterations is raised, runtime or memory consumption can increase notably and might result in runtime or memory issues so that method failure is, after all, not solved.
If, after these considerations, a change in the implementation of the given method is deemed reasonable, any deviation from the default parameter setting of a given method should be reported and justified transparently. Note that, as will be discussed in Section 3.3.5, changes in a method’s implementation could also be applied as a fallback strategy.
3.3.5 Consider a fallback strategy
When contemplating ways to mimic real-world users when encountering method failure in later research stages, Morris et al., (2019) argue that they would usually not give up but naturally resort to an alternative method, a fallback, that successfully provides an output. Therefore, it can make sense to transfer this fallback strategy into the context of methodological comparison studies and evaluate the performance of method pipelines of the type
“use method A if it successfully produces an output, use method B otherwise”.
On the one hand, this approach incorporates what data analysts really do in practice, and on the other, it indirectly yields automatic handling of undefined performance values in case they affect method A but not method B. If chosen carefully and in a way that there is a suitable fallback for each instance of method failure, fallback strategies allow for the assessment of (absolute and relative) unconditional performance of the methods under investigation. Besides realistically reflecting the behavior of real-world users, this approach results in a performance assessment that depends only on the definition of the pipeline. Note that when using fallback strategies, we no longer evaluate the “pure” performance of method A. However, it is clear from Section 3.1 that this in an unachievable goal in the first place.
The concept of fallback strategies is already intuitively considered in some existing comparison studies. Voeltz, (2024), for instance, conclude from their study that “the approach by Frömke et al., (2022) could be used as a fallback strategy in case of non-convergence [of the overall best-performing method] because it always yields results”, where the approach by Frömke et al., (2022) is a non-iterative procedure, therefore never affected by non-convergence by design. Going one step further than Voeltz, (2024), one could provide a performance assessment for the resulting pipeline.
A special case of the fallback strategy has indirectly been addressed in Section 3.3.4: it may sometimes make sense to define an alternative implementation of the given method (e.g, using a different optimizer or an increased number of iterations) as a fallback instead of a completely different method (as in Crowther et al., (2014) who changed the number of quadrature points). The corresponding pipeline would then consist of the default implementation of that method for those data sets where it successfully produces an output and the alternative implementation for the remaining data sets.
In practice, the choice of a suitable fallback depends on the research phase. For instance, in early phases, the fallback may be constituted of an alternative implementation (e.g., alternative parameter choice) of the original method. Also, authors should strongly base the selection of fallbacks on what can be considered realistic given the expertise in statistics and programming of the targeted users. For instance, in a study targeting data analysts with limited statistical background, the fallback strategy may be an alternative method with default parameters rather than the same method with complex parameter settings. Since the choice of suitable fallbacks involves some subjectivity, authors should justify their choice(s) transparently.
If failures are expected at the planning stage, i.e., before conducting the study, one may consider the whole pipeline as a competing method to be assessed in the comparison study. Another approach would be to design a pipeline after seeing the results of the comparison study, as did by Voeltz, (2024). In this case, the definition of the pipeline can be viewed as a result of the comparison study rather than a method that has to be evaluated in the first place. The definition of the fallback strategy, however, may not always be as straightforward as in Voeltz, (2024). It may for example make sense to consider complex rules involving a second fallback in case the first one fails.
3.3.6 Transparently report failure proportions, underlying interplays, and applied approaches
Following all recommendations above, authors should transparently report (i) any occurrences and proportions of method failure, (ii) the underlying interplay(s) causing failure (in different instances), and (iii) the corresponding approaches following the in-depth investigations of method failure.
Report any occurrences and proportions of method failure  This can serve as an important criterion for real-world users selecting a suitable method. Even the “best performing” method is of limited use if the amount of settings in which it can be applied without frequent failures is limited. Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Voeltz, (2024) for an example. If method failure is predictable for certain settings and all affected methods fail for the same data sets (see, e.g., the example from Section 4.1), it may make sense to include the binary performance measure of the type “can handle setting: yes. vs. no” and report it alongside the corresponding number of repetitions/benchmark data sets associated with the failure.
Report interplay of method and data causing failure  Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”. This enables a more comprehensive overview of the applicability of the methods compared to reporting the corresponding failure proportions only. Readers can learn in which settings a method is applicable, and no less importantly, in which it is not. Additionally, these insights serve as justification of the chosen handlings to method failure. An example of a corresponding reporting can be found in Hornung et al., (2024) who provide detailed descriptions in the supplement of their paper.
Report applied handlings  Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail. This includes the removal of any misused methods, the removal of data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and the incorporations of fallback strategies. Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at), and the research stage in which it takes place.
3.3.7 Finally: think ahead, pre-register but remain flexible
The previous examples and those that will be outlined in Section 4 make clear that method failure and its cause(s), related to complex dynamics between factors such as method and data set, are often difficult to anticipate. While pre-specifying the handling of undefined values has the advantage of reducing analytical uncertainty and its dangers in terms of selective reporting (Siepe et al., , 2024), pre-specifying a standard procedure prior to conducting the study may sometimes be delicate.
Our recommendation is, as far as possible, to consider in advance which combinations of data and methods could lead to failure. For instance, carefully reading user manuals and help pages of all methods before conducting the study might reveal specific requirements of a method on the data. This may make some instances of method failure less unexpected. More generally, making method failures less unexpected can be achieved by conducting a pilot experiment. Pilot experiments test whether an experimental design is appropriate for the specific context in which it is applied, and are recommended in several textbooks on Design and Analysis of Experiments (Cohen, , 1995; Dean et al., , 2017; Montgomery, , 2020). For example, in the context of real data comparison studies, a pilot experiment might use a fraction of the available data sets, say $20\%$, to test whether the specified method implementations produce reasonable results. The remaining $80\%$ of the data sets are then used for the main experiment, which may be refined based on what is learned in the pilot. A pilot study can also help authors anticipate possible runtime issues for which they can then pre-specify the random subset of data sets on which the experiment for the affected method will be restricted. Of course, it is important to clearly report the results of the pilot study too. After implementing these strategies, pre-specifying a procedure for handling method failure may be less hazardous. However, even despite pre-registration, researchers should be aware that unforeseen issues may appear and require adaptations of the pre-specified strategy. See Lakens, (2024) for a discussion of the consequences of deviations from pre-registration on the severity and validity of inferences.
3.4 Slimmed-down recommendations when time is short
For some authors, there may be time constraints that do not allow for any (in-depth) implementations of our recommendations from the previous Section 3.3, however, authors nevertheless wish to aggregate results over data sets despite the existence of undefined values caused by method failure. In these cases, where it is difficult to follow the entirety of our recommendations, we recommend the following slimmed-down analysis and reporting strategy. Authors may present performance results after (i) discarding the corresponding data sets for the failing methods, (ii) discarding these data sets for all methods, and (iii) report the failure proportions for all methods, ideally as an additional performance measure. This three-fold analysis and reporting enables more differentiated assessments of the methods’ performances and avoids the preference of methods based solely on how data sets associated with method failure are discarded. In particular, it provides insights into how the methods successfully producing an output perform on the data sets for which others fail (and how this can influence the method comparison).
However, following Section 3.1, authors should be aware and clearly state in their manuscript that any form of aggregating performance results despite existing undefined values does not allow an assessment of unconditional performance of all methods under investigation.
4 Illustrations
In the following, we illustrate the guidelines presented in Section 3 through two illustrative comparison studies featuring method failure: a statistical simulation study and a benchmark study at the interface between statistics and predictive modeling. For both examples, we first demonstrate how the ad hoc approaches discussed in Section 3.2, such as discarding data sets associated with failure, can distort the results and even lead to misleading method recommendations. Then, we illustrate the implementation of the recommendations from Section 3.3. Thereby, our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison, rather than the method comparison itself. This has three consequences. First, both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study. Second, we do not give a comprehensive overview of all aspects of the studies but focus on those that are relevant to the issue of undefined performance values. Third, our goal is not to provide insights into the performance of the compared methods (larger dedicated studies would be necessary for this purpose).
4.1 Example 1: Comparing methods for the estimation of odds ratios
4.1.1 General
Our first example, a statistical simulation study, deals with the comparison of odds ratio estimation methods. Odds ratios (“ORs”) quantify the association between an exposure and an outcome and are typically used in (retrospective) case-control studies. Here, we focus on the case of a binary exposure $X\in\left\{0,1\right\}$ and binary outcome $Y\in\left\{0,1\right\}$. While a straightforward way of estimating the OR is to calculate $\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$ (where $n_{ij}$ is the observed number of subjects with $X=i$ and $Y=j$), other methods employ more complex modeling strategies. Note that ORs are bounded in $(0,\infty)$, where values $>1$ imply a positive association between exposure and outcome.
With the true OR value being generally unknown in real data sets, OR estimation methods are typically compared using simulation studies under different scenarios, e.g., different sample sizes, underlying true OR values, and probabilities of exposure $p_{x}$. Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations. This results in $n_{ij}=0$ for the corresponding combination $i,j\in\{0,1\}$ for this data set. These zeros are called “sampling zeros” and cause computational issues (e.g., division by 00) for many of the available OR estimation methods. They lead to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$”, resulting in undefined performance values in the corresponding data sets. In practice, users often apply the “Haldane-Anscombe correction” (Haldane, , 1956; Anscombe, , 1956) to handle sampling zeros in the OR estimation. This correction consists of adding an offset of $+0.5$ to all four cells in the contingency table to avoid issues in computation.
4.1.2 Design of the fictive study
In our fictive simulation study, we vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size at the fixed value of $n.obs=50$. This results in eight distinct simulation scenarios (see Table 5 for an overview). For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets (where the exposure variable $X$ is generated using a Bernoulli distribution with probability $P(X=1)=p_{x}$ and outcome $Y$ is generated using the logistic function with the linear effect of exposure $X$ specified as $\beta_{X}=\log(\text{OR})$). Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR, namely “Fisher”, “Midp”, “Small” (all three implemented in R package epitools (Aragon et al., , 2020)), “Woolf” (Woolf et al., , 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard, , 2019)), and “Manual”, which corresponds to the manual computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$. Note that we omit method “Wald” from the epitools package since its estimation output is identical to the manual computation (i.e., method “Manual”). From the five estimators under investigation, only Small and Woolf can handle sampling zeros internally. For each OR estimator and simulation scenario, bias on the $\log$-scale is computed as
[…]
and the MSE on the $\log$-scale is obtained as
[…]
Note that the $\log$-transformation ensures symmetry around the value of “no association”. For both performance measures, values closer to 00 indicate better performance (Morris et al., , 2019).
4.1.3 Occurrence and handling of method failure
Table 5 displays the proportion of the 100 000100000100\,000100 000 simulated data sets containing at least one sampling zero across the eight simulation scenarios. As expected, this proportion increases with an increasing true OR$>1$, while decreasing when exposure $X$ is more balanced.
Table 5: Overview of the proportion of the 100 000100000100\,000100 000 simulated data sets with sampling zeros across the simulation scenarios. $n.obs$: number of observations, $p_{x}$: exposure probability.
Simulation scenario n.obs true OR $p_{x}$ Sampling zero proportion
1 50 2 $0.25$ 1.28%
2 50 3 $0.25$ 3.96%
3 50 4 $0.25$ 7.8%
4 50 5 $0.25$ 12.0%
5 50 2 $0.5$ 0.01%
6 50 3 $0.5$ 0.11%
7 50 4 $0.5$ 0.50%
8 50 5 $0.5$ 1.27%
Our analysis consists of two parts. First, we compare the performance of the OR estimation methods regarding performance measures bias and MSE when ignoring data sets with sampling zeros for all methods versus ignoring data sets only for the failing methods (“ad hoc approaches”). While we are aware that users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.3.5. As fallbacks for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf. For the manual estimation of the OR (method “Manual”), we only apply the Haldane-Anscombe correction as a fallback, assuming that a user applying manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort. Together with Woolf and Small, nine “methods” (“pure” methods or pipelines) are compared in total and the results are presented in Section 4.1.4.
Note that since we focus on the effect of handling method failure (instead of the actual results of the comparison), and this effect becomes clearer through examining the individual scenarios, we do not aggregate the performances across the eight scenarios in either of the two analysis parts.
Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios, derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B). In both subfigures, the ranks are compared between discarding data sets yielding failure for the failing methods only (denoted by “Failing” on the $x$-axis) versus for all methods (denoted by “All” on the $x$-axis). A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario. Note that Small and Woolf are the only methods in this selection that can deal with sampling zeros internally. Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR).
Ad hoc approaches  For the first analysis, Figure 1 shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison. This particularly affects estimator Woolf, one of the two methods that can handle sampling zeros internally. In terms of bias (Figure 1 (A)), its performance rank deteriorates in most scenarios and by up to three positions when discarding data sets with sampling zeros for all methods (including Woolf and Small). The reason for this is that across the simulated data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it. This leads to top performance when including all data sets in its performance assessment (while discarding data sets with sampling zeros for the failing methods). Even more so, the manner of discarding data sets with sampling zeros affects Woolf’s performance rank regarding MSE (Figure 1 (B)). In some scenarios, Woolf’s rank improves from bottom to top rank, indicating its poorer performance for data sets with sampling zeros. The performance rank of Small, the second method that can handle sampling zeros internally, is less affected by this for both bias and MSE. In particular, it scores last in terms of bias across all scenarios and for both manners of discarding (Figure 1 (A)). However, its performance rank changes by up to two positions for MSE (Figure 1 (B)). The performance ranks of the remaining methods are also affected regarding bias and MSE, where the manner of discarding data sets with sampling zeros affects their performance ranks by up to two positions.
Note that the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros. In particular, scenarios 5 and 6, with a maximum of $0.11\%$ of data sets containing sampling zeros, yield equal performance rankings across both handlings. However, differences in rankings within a given scenario are already observable with a proportion of $0.5\%$ of data sets with sampling zeros.
Figure 2: Performance ranks of the OR estimation methods across the eight simulation scenarios when applying fallback strategies to methods Midp (“Mid”), Fisher (“Fis”), and Manual (“Man”). The rankings are derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B). A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario. Note that for Midp and Small, the fallbacks Small (“Sma”), Woolf (“Woo”), and the Haldane-Anscombe correction (“$+0.5$”) are used. For Manual, only the Haldane-Anscombe correction is used as fallback. Small and Woolf are the only methods in this selection that can deal with sampling zeros internally, therefore not requiring any fallback. This amounts to nine “methods” in total. The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$”.
4.1.4 Recommended handling
As expected, the relative performance between Woolf and Small remains the same between applying fallbacks for the other methods (Figure 2) versus discarding data sets with sampling zeros for the failing methods only (“failing” parts in Figure 1 (A) and (B)). This makes sense since neither method requires a fallback, resulting in unchanged bias and MSE. Midp’s overall good performance regarding bias and MSE is underpinned by using fallbacks Small and the Haldane-Anscombe correction. Here, it is debatable which fallback is preferable. Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios. However, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp. Note the interesting observation that Small scores below the remaining methods regarding bias when applied alone, but performs well as a fallback for Midp.
4.2 Example 2: Comparing methods for computing confidence intervals for the generalization error
4.2.1 General
The fictive study aims to compare the performance of two methods for constructing confidence intervals (CIs) for the generalization performance of a prediction model based on resampling-based estimates, such as cross-validation. The “naive” method “$N$” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “$C$” corrects for it. The performance of both methods is assessed in terms of coverage, which is the probability that a constructed CI includes the true value of the generalization performance. It should be close to its nominal value, e.g. $0.95$ if we consider $95\%$-CIs.
4.2.2 Design of the fictive study
Classification trees (Breiman et al., , 1984) are considered as models for predictive modeling. As benchmark data set, we use the data set provided by the SAPA project (Condon et al., , 2017), further processed according to Klau et al., (2023). The response variable is binary and five variables from the data set are used as predictors. The generalization performance is quantified using the Area Under the Curve (AUC), where a value of $0.5$ corresponds to a random prediction and a value of $1$ indicates perfect classification. The AUC is estimated using $15$ repetitions of repeated subsampling ($i=1,\ldots,15$) with a 4:1 ratio (i.e., the data is randomly split $15$ times into $80\%$ on which the model is trained and $20\%$ on which the AUC is obtained (Hastie et al., , 2009)).
Method $C$ suggested by Nadeau and Bengio, (2003) computes the CI for the generalization AUC as
$\Big{[}\ \overline{\widehat{AUC}}-t_{0.975,14}\cdot\sqrt{\Big{(}\frac{1}{15}+c\Big{)}\cdot S^{2}_{\widehat{AUC}_{i}}},\quad\overline{\widehat{AUC}}+t_{0.975,14}\cdot\sqrt{\Big{(}\frac{1}{15}+c\Big{)}\cdot S^{2}_{\widehat{AUC}_{i}}}\ \Big{]},$ (1)
where $\overline{\widehat{AUC}}$ represents the mean AUC estimate over the $15$ subsampling repetitions and $S^{2}_{\widehat{AUC}_{i}}$ the corresponding sample variance, while $t_{0.975,14}$ is the $95\%$-quantile of the $t$-distribution with $15-1=14$ degrees of freedom. The fixed correction term $c>0$ incorporates the correlation structure between the subsampling repetitions (Nadeau and Bengio, , 2003). Method $N$, on the other hand, does not apply any correction, i.e. sets $c=0$ in the above formula.
[…]
The study is conducted using R package mlr3, method $N$ is implemented using function t.test() from package base and for method $C$, $\overline{\widehat{AUC}}$ and $S^{2}_{\widehat{AUC}_{i}}$ are obtained using functions mean() and sd() (from the same package).
Figure 3: Design of the fictive study to obtain the empirical coverage of methods $C$ and $N$. An approximation of the “true AUC” is obtained based on the very large $D_{test}$.
4.2.3 Occurrence and handling of method failure
For method $N$, the implementation of the above-described procedure returns the following error message for 300 of the $1000$ iterations in R:
Error in t.test.default: data are essentially constant.
Hence, the confidence intervals cannot be constructed for method $N$ in these iterations, making it impossible to assess whether they cover the true AUC. Method $C$ is not affected by these errors. At first glance, the error message gives little indication of how to resolve it. In the following, we first apply three ad hoc approaches to handle these errors. Then, by following our in-depth guidelines from Section 3, we demonstrate how these ad hoc approaches prove to be inadequate, and illustrate more suitable ways to handle method failure in this example.
Ad hoc approaches  One could be inclined to apply one of the following ad hoc approaches:
  • 1.
    for method $N$, ignore all data sets in which it fails (method $C$ is unaffected).
  • 2.
    for both methods, ignore those iterations in which $N$ fails.
  • 3.
    for method $N$, set “undefined” confidence intervals to “not covering the true AUC”.
Note that the intuition to approach 3, which can be viewed as a form of imputation, is that a CI that does not exist cannot cover any value. It thus induces a negative correlation between the proportion of undefined confidence intervals and the coverage of method $N$.
Table 6: Empirical coverages of methods $N$ and $C$ resulting for (a) the three “ad hoc” approaches and (b) setting undefined CIs to zero-width CIs.
(a) Ad hoc approaches: NA-CIs are (b) NA-CIs are
discarded for
method $N$ only
discarded for
methods $N$ and $C$
set to “not covering
true AUC”
set to zero-width
interval
Method $N$ $0.54$ $0.54$ $0.38$ $0.68$
Method $C$ $1$ $1$ $1$ $1$
The coverages resulting from the ad hoc approaches are displayed in Table 6 (a). While the coverage of method $C$ is constant across all three ad hoc approaches (coverages above the nominal level indicate conservativeness), method $N$, while too liberal across all approaches, is even more liberal using the ad hoc imputation approach.
4.2.4 Recommended handling
We now closely inspect the modeling processes underlying the study. The goal is to find out the root cause of the error message and to adapt the handling of the undefined performance values accordingly. The errors in method $N$ occur when the estimated AUCs are equal across all $15$ subsampling repetitions, resulting in a sample variance $S^{2}_{\widehat{AUC}{i}}$ of 0. This zero variance in the denominator of the $t$-statistic causes an error in function t.test(). However, focusing on the confidence interval, we realize that from a theoretical point of view, a sample variance of 0 simply means that the CI is the zero-width interval $\left\{\overline{\widehat{AUC}}\right\}$. Therefore, method $N$ not yielding a CI (and returning an error) is a technical artifact of its implementation. A straightforward solution would be to modify the implementation of method $N$, using mean() and sd() instead of t.test(), so that it computes the zero-width CI, which can enter the calculation of coverage as any other CI (see Table 6 (b) for the resulting coverage).
However, taking the investigation another step further reveals that a sample variance $S^{2}_{\widehat{AUC}{i}}=0$ always occurs alongside all $15$ estimated (and the approximate true) AUC being equal to $0.5$. Ultimately, this worst possible performance of the classification tree (“random prediction”) occurs when the classification tree does not perform any splitting. This is related to the characteristics of the data set. For example, the problem disappears if we include additional predictors: splitting then occurs, leading to an AUC higher than $0.5$.
The adequate handling of method failure in this illustrative study, as described in Section 3, depends on the aim of the study. In particular, a prediction problem in which splitting often does not take place can be viewed as an extreme scenario. If the study’s scope includes such extreme scenarios, one could attempt to suitably modify parameters of the classification tree to favor splitting. If, however, the study aims to compare both methods in regular (i.e., less extreme) scenarios, one might consider additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
Finally, note that this example illustrates that method failure may be complex and that the interplay causing it may not be immediately identifiable, even if the benchmark study consists of a single scenario only. In a real benchmark study which typically contains multiple scenarios, thorough investigations can become even more difficult.
5 Discussion
Method failure in comparison studies, often manifesting as issues regarding calculation (e.g., non-convergence), runtime, or memory issues, complicates the desired comparison of unconditional performance between the methods under investigation, i.e. across all simulated or benchmark data sets. Our informal literature review and illustrations both make clear that authors should not blindly rely on the initially perceived manifestation of failure but rather make extensive investigations of the sources of failure, which often guides the choice of the handling of the resulting undefined values.
Even if we argue for decisions on a case-by-case basis after careful inspection of the failures, we show and illustrate that some general principles hold in most situations. In particular, method failure usually leads to non-existing values instead of just “missing”, i.e., existent but unobserved values. This means imputation, a common but missing-data-inspired approach to handling method failure, is often inadequate. Another popular and seemingly logical approach, namely discarding data associated with failure (either for all or failing methods), almost never enables the comparison of unconditional performance. In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
An approach that in our view deserves more attention than it receives currently in the literature consists in evaluating whole “pipelines” including one or several fallback strategies in case of method failure. This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one. Future work is required to collect experiences and define standards regarding such pipelines and their evaluation. Above all, transparent reporting of the frequency of failure and of the corresponding handling is always recommended.
A limitation of our recommendations is that, beyond obvious cases, they typically require both high time resources and thorough expertise in the underlying modelling processes and methods. Researchers conducting comparison studies may not always have these resources to the full extent. This particularly applies to a nuanced examination of the interactions between methods and data sets leading to method failure in different instances. We understand that this bottleneck may impair the full implementation of our recommendations in practice, hence our attempt to provide a slimmed-down version of the recommendations. When time resources and expertise are restricted, transparent reporting is all the more important. Also, our viewpoint on method failure as leading to “undefined” performance values does not always hold unrestrictedly, which can complicate the implementation of our recommendations. However, with method failure being a complex matter in and of itself, this underscores the importance of a case-by-case evaluation, as described above.
In this article, method failure has been treated as black-or-white. However, in practice, whether or not a method has failed is itself a value of judgment. One such example is near-separation (van Smeden et al., , 2016), which does not necessarily make itself obvious, hence the existence of dedicated methods to detect it (Mansournia et al., , 2017). Such grey-area failures add further complexities to benchmark and simulation studies: overall performance may depend on the specific criteria used to judge failure.
To sum up, we believe there is not “one” right way to handle method failure in all situations. Instead, it is important to carefully choose an approach depending on the context, and then report, explain (and ideally discuss) the decisions made. With our work, we hope to contribute meaningful ideas on how to think about method failure. They lead to helpful strategies for dealing with it, as outlined as part of our recommendations. However, we also believe that undefined performance values are only one of many issues related to comparison studies that deserve much more time, efforts and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020). We need more well-designed neutral comparison studies (Boulesteix et al., , 2017)—late phase studies according to the terminology of Heinze et al., (2024). In such studies, no time is spent on the development of new methods. This leaves time for implementing our recommendations and diving into the technical details of the methods under investigation.
Funding Information
This work is supported in part by funds from the German Research Foundation (DFG: BO3139/7 and BO3139/9-1). TPM was funded by the UK Medical Research Council (grant MC_UU_00004/09).
Acknowledgements
The authors thank F. Julian D. Lange for useful literature input and Clemens Kreutz for valuable comments. The authors further thank Savanna Ratky for language corrections and Luzia Hanßum for support with formatting the manuscript.
Conflicting interests
The authors have declared no conflicts of interest for this article.
Data availability statement
The code to reproduce all results from Section 4 are available on Github.